# An Empirical Study of Cross-Lingual Transfer Learning in Programming Languages

### v2

@@ IV-A Transfer analysis
We explore RQ1 via extensive experiments on every combination of source and target programming language, for each task. The four heat maps of Fig. 5 show the results. A source programming language is the language used for fine-tuning a model, with samples in the training data. A target programming language is the language used to evaluate a model, with samples in the test data. Fig. 5 shows source languages on the X-axis and target languages on the Y-axis. Monolingual scores (where the source and the target language are the same) are highlighted with bold boxes.
For RQ1 we analyze model performance trends with respect to task, source language, and target language. We provide some practical insights, like which languages are the toughest target languages and which are the best source languages. We also discuss some interesting source-and-target language pairs.
Task dependency
@@
The heat maps in Fig. 5 show that transferability of source languages varies depending on the task. Visually this can be seen by how the color gradient varies for each task. To quantify this observation, we calculate the mean scores across all the language combinations. The mean cross-lingual scores are 0.78 for clone detection, 0.75 for code repair, 0.65 for error detection, and 0.47 for solution domain classification. The monolingual scores (in bold boxes) are higher (more red) than the cross-lingual scores, averaging 0.91 for clone detection, 0.98 for code repair, 0.78 for error detection, and 0.51 for solution domain classification. The zero-shot columns, where the model is not finetuned on any source language, have lower scores (more blue) for all heat maps. The mean zero-shot scores are 0.49 for clone detection, 0.28 for code repair, 0.49 for error detection, and 0.01 for solution domain classification. Variations in the mean monolingual, cross-lingual, and zero-shot scores across tasks confirm that model performance is task dependent. The difference in overall zero-shot and cross-lingual scores indicates that in the absence of fine-tuning data for any task in any programming language, it is better to finetune with some other language in that task than using a zero-shot setting. This cross-lingual advantage over the zero-shot setting occurs for every task.
@@
Looking at these scores in absolute terms reveals that the choice of tasks also affects transferability. For solution domain classification, the difference in mean cross-lingual and monolingual scores is only 0.04, whereas for code repair, it is 0.23. The difference in mean cross-lingual and zero-shot scores for the same two tasks is 0.46 and 0.48. Although overall trends show learning transfers well in all tasks, cross-lingual performance in relation to monolingual performance is highly task dependent. Within each task, model performance depends on the choice of source and target programming language, which we explore next. Tab. II in the supplementary material shows the mean score by task.
Target language dependency
@@
Target programming languages fall into two categories, high- and low-resource. In Fig. 5, high-resource target programming languages are represented on both axes. Low-resource target programming languages are represented only on the Y-axis. Models are finetuned only on high-resource programming languages and tested on both high- and low-resource programming languages.
@@
To understand overall trends across all four tasks, we consider ten target programming languages that are common to all four tasks. Six (Kotlin, JavaScript, Java, Python, C, and C++) are high-resource in all four tasks. Four (Go, Rust, PHP, and C#) are low-resource in error detection. For each of these ten languages, for each of the four tasks, we calculate the mean cross-lingual score using all scores in the row of Fig. 5 for each language, excluding the monolingual score in the bold box and the zero-shot score in the extreme left. We exclude monolingual and zero-shot scores so as to understand transferability from different languages to the target language. Based on this mean we rank each of the 10 languages for each task, and then calculate the mean rank for each language.
We find Java, Go, Rust, JavaScript, and Kotlin are target languages that benefit most from transfer learning, while PHP, C#, C, C++, and Python benefit least, with Java, Go being the best and Python, C++ being the worst. C++ has a reputation as being one of the hardest programming languages to learn for human software developers, which may explain our finding that it is a difficult target for LLM transfer learning. Some target programming languages are good at monolingual performance, but much worse at cross-lingual performance; one example across multiple tasks is C++. In-language performance for C++ is 0.95, 0.99, and 0.46 for clone detection, code repair, and solution domain classification, whereas its mean cross-language performance is 0.79, 0.81, and 0.41 for the same tasks. Java is an example of more consistent in-language and cross-language transferability across tasks and hence can be considered relatively more transferable than C++. Java has in-language scores of 0.95, 0.99, and 0.45 for clone detection, code repair, and solution domain classification, and cross-language scores of 0.86, 0.84, and 0.42 for the same. All numbers are reported Tab. III in the supplementary material.
@@
We do a similar analysis for low-resource target languages in clone detection and code repair, since these two tasks have a wider and more varied set of low-resource languages. Here we find Dart, TypeScript, Elixir, Lua, and Swift benefit most from cross-lingual training, with Dart being the best. Dart, TypeScript have average scores of 0.87 and 0.79, 0.82 and 0.83 for clone detection and code repair. COBOL, F#, Octave, bc, and Vim are least likely to benefit, with Vim being the worst. Vim has average scores of 0.69, 0.67 and COBOL has average scores of 0.72, 0.74 for Clone Detection and Code Repair. Tab. IV in the supplementary material has more information on mean rank and average score of low-resource target languages for clone detection and code repair.
@@
For error detection, only a smaller set of languages can be considered low resource. We observe low variation in average scores across different sources with highest being Rust with 0.66 and lowest being C# with 0.64. For solution domain classification there are no low-resource target languages.
Our analysis shows that transferability depends on the choice of task and target language. We can see by comparing with zero-shot scores that all target programming languages benefit from cross-lingual training, however some languages benefit much more than others.
Most Transferable Source Language
@@
The number of source languages vary for each of the four tasks. To understand the source language dependency and to identify the most transferrable source language, we start from languages all four tasks have in common (JavaScript, C, Kotlin, Java, C++, and Python). For each task, for each language, we calulate the mean cross-lingual score. As an example, consider the column for source language Python on the X-axis of Fig. 5(a). We calculate the mean over the scores except 0.94, which is the monolingual score with target language Python. Once we have the cross-lingual mean for each language, we rank the source languages based on their mean score for each task. Based on these rankings, we calculate the mean rank, for overall cross-lingual performance of source languages, across all target languages, across all tasks.
@@
The most transferrable source languages from best to worst are Kotlin, JavaScript, Java, Python, C, and C++. It may be surprising that Kotlin is the best-performing source language, since it is rarely discussed in the context of training LLMs and is not part of the pre-training corpus for CodeT5 [22]. Prior work on transfer learning for natural languages [17] yielded a similar surprise, with Romanian as the best-performing source language. A possible explanation could be that Romanian and Kotlin both have strong roots in dominant languages of the past (Latin and Java), while sitting at a cross-roads that helped them absorb many influences. Detailed scores and rankings are in Tab. V in the supplementary material.
@@
The above analysis considers all target languages, including high-resource target languages. In a more realistic setting, we can ignore high-resource target languages because the monolingual option is available for every high-resource language in every task. If we repeat the above analysis with only low-resource languages, we are left with a smaller set of target languages and only three tasks (all target languages for solution domain classification are high-resource). For the remaining three tasks and six common source languages, analysis shows that the most transferrable source languages are JavaScript, Java, Kotlin, C, Python, and C++, with JavaScript being the best and C++ being the worst. JavaScript and Java contribute the most data to CodeT5 pretraining; Kotlin and C++ are absent [22]. For details see Tab. VI in the supplementary material.
@@
If we consider only clone detection and code repair, we get a much broader set of about 20 common source languages. Repeating the analysis for these two tasks for low-resource target languages only, we find once again that JavaScript is the most transferable and C++ is among the least. In this case the most transferable languages are JavaScript, Scala, and Haskell, while the least transferable are Fortran, AWK, C++, Lisp, and Pascal. More details are in Tab. VII in the supplementary material.
@@
For clone detection in Fig. 5(a), notice that a model finetuned on Ruby does almost as well on Crystal, a language whose syntax is inspired by Ruby. Fig. 5(a-b) also shows that surprisingly, Haskell does very well on COBOL, much better than FORTRAN does on COBOL. From the same heat maps notice Lisp does very well on Scheme, which can be expected as Scheme is a dialect of Lisp. Similary Python does very well on Cython. Fig. 5(a-d) also show that C-category languages (C, C++, and C#) do well on each other. Details are in Tab. VIII in the supplementary material.
Dendrograms
Dendrograms for each task cluster programming languages with similar transfer learning performances. In the code repair task, syntaxically and semantically related languages are next to each other. For example, C is paired with C++ and is next to the pair C# and Java. JavaScript is paired with TypeScript, Python with Cython, Lua with Moon, Lisp with Scheme. On the other hand, relatedness of languages plays a smaller role in the dendrograms for other tasks. This might be because while the code repair task requires deep language understanding, other tasks depend more upon other features. For instance, Ahmed and Devanbu [11] have shown that for some tasks, identifiers are more important. Section IV-C will analyze the relationship between features of language pairs and their transfer performance.
1. Learning transfers well for all tasks. Cross-lingual learning transfers better than zero-shot. 2. Learning depends on the target language. Among low resource languages, Dart and TypeScript are the best while Vim and COBOL are among the worst. 3. Kotlin and JavaScript are among the best target languages, C++ is among the worst. 4. Some cross-lingual language pairs show particularly good results, such as Ruby and Crystal, Haskell and COBOL.
IV-B Performance Prediction

### v3

@@ 4.1 Transfer analysis
We explore RQ1 via extensive experiments on every combination of source and target programming language, for each task. The four heat maps of Figure 5 show the results. A source programming language is the language used for fine-tuning a model, with samples in the training data. A target programming language is the language used to evaluate a model, with samples in the test data. Figure 5 shows source languages on the horizontal axis and target languages on the vertical axis. Monolingual scores (where the source and the target language are the same) are highlighted with bold boxes. The first column from the left indicates zero-shot performance of the base Code-T5 model on different target languages. For RQ1 we analyze model performance trends with respect to task, source language, and target language.
Task dependency.
@@
The heat maps in Figure 5 show that transferability of source languages varies depending on the task. Visually this can be seen by how the color gradient varies for each task. To quantify this observation, we calculate the mean scores across all the language combinations. The mean cross-lingual scores are 0.78 for clone detection, 0.75 for code repair, 0.65 for error detection, and 0.47 for solution domain classification. The monolingual scores (in bold boxes) are higher (more red) than the cross-lingual scores, averaging 0.91, 0.98, 0.78, and 0.51 for the same four tasks. The zero-shot columns, where the model is not finetuned on any source language, have lower scores (more blue) for all heat maps, averaging 0.49, 0.28, 0.49, and 0.01. Variations in the mean monolingual, cross-lingual, and zero-shot scores across tasks confirm that model performance is task dependent. The difference in overall zero-shot and cross-lingual scores indicates that in the absence of fine-tuning data for any task in any programming language, it is better to finetune with some other language in that task than to use a zero-shot setting. This cross-lingual advantage holds for every task.
Mean Score Task Mono Cross Overall Zero-Shot Clone Detection 0.91 0.78 0.79 0.49 Code Repair 0.98 0.75 0.76 0.28 Error Detection 0.78 0.65 0.66 0.49 Solution Domain Classification 0.51 0.47 0.47 0.01
Table 2: Mean Scores by Task. Monolingual implies finetuning where the train and test data language is the same. Cross-lingual implies finetuning where the train and test data languages differ. Overall scores include both monolingual and cross-lingual scenarios. Zero-shot means the performance of the base pre-trained model (CodeT5-220M) on the test set without finetuning.
@@
Looking at these scores in absolute terms reveals that the choice of tasks also affects transferability. For solution domain classification, the difference in mean cross-lingual and monolingual scores is only 0.04, whereas for code repair, it is 0.23. The difference in mean cross-lingual and zero-shot scores for the same two tasks is 0.46 and 0.48. Although overall trends show learning transfers well in all tasks, cross-lingual performance in relation to monolingual performance is highly task dependent. Within each task, model performance depends on the choice of source and target programming language, which we explore next. Table 2 shows the mean score by task.
Learning transfers well for all tasks. Cross-lingual learning works better than zero-shot.
Target language dependency.
Clone Detection Code Repair Error Detection Solution Domain Target cross 0-shot cross 0-shot cross 0-shot cross 0-shot Mean Rank Java 0.86 0.44 0.84 0.25 0.66 0.60 0.42 0.009 4.25 Go 0.86 0.50 0.71 0.16 0.66 0.55 0.51 0.003 4.25 Rust 0.80 0.49 0.77 0.23 0.66 0.36 0.53 0.001 4.50 JavaScript 0.83 0.48 0.85 0.34 0.65 0.49 0.52 0.008 4.75 Kotlin 0.84 0.49 0.76 0.17 0.66 0.48 0.50 0.005 5.00 PHP 0.84 0.51 0.82 0.28 0.65 0.38 0.50 0.004 5.25 C# 0.84 0.48 0.82 0.2 0.64 0.54 0.47 0.046 5.50 C 0.83 0.45 0.82 0.34 0.66 0.57 0.48 0.004 5.75 C++ 0.79 0.46 0.81 0.24 0.66 0.51 0.41 0.004 7.00 Python 0.83 0.50 0.76 0.20 0.64 0.49 0.37 0.006 8.75
Table 3: Score distribution by target languages common for all tasks. Scores are the mean of the score of every source for a given target language. Mean rank is used to rank the languages based on their ranking for each task. All the languages shown are high-resource target languages for Clone Detection, Code Repair and Solution Domain Classification, which means they are also among the source languages for those tasks. C#, Go, PHP, Ruby, and Rust are low-resource languages for Error Detection. The table shows that all languages benefit from transfer learning. Java and Go seem to benefit the most and C++ and Python seem to benefit the least.
@@
In Figure 5, high-resource target programming languages are represented on both axes. Low-resource target programming languages are represented only on the vertical axis. Models are finetuned only on high-resource programming languages and tested on both high- and low-resource programming languages.
@@
To understand overall trends across all four tasks, we consider ten target programming languages that are common to all four tasks. Six (Kotlin, JavaScript, Java, Python, C, and C++) are high-resource in all four tasks. Four (Go, Rust, PHP, and C#) are low-resource in error detection. For each of these ten languages, for each of the four tasks, we calculate the mean cross-lingual score using all scores in the row of Figure 5 for each language, excluding the monolingual score in the bold box and the zero-shot score in the extreme left. Based on this mean we rank each of the 10 languages for each task, and then calculate the mean rank for each language. The most transferrable target languages from best to worst are Java, Go, Rust, JavaScript, Kotlin, PHP, C#, C, C++, and Python. Some target programming languages are good at monolingual performance, but much worse at cross-lingual performance. For example, in-language performance for C++ is 0.95, 0.99, and 0.46 for clone detection, code repair, and solution domain classification, whereas its mean cross-language performance is 0.79, 0.81, and 0.41 for the same tasks. This might be explained by the fact that C++ has a reputation as being one of the hardest programming languages to learn for human software developers. All numbers are in Table 3.
Clone Detection Code Repair Target Language cross 0-shot cross 0-shot Mean Rank Dart 0.87 0.51 0.79 0.28 2.0 TypeScript 0.83 0.49 0.83 0.25 3.5 Elixir 0.83 0.51 0.75 0.21 5.0 Lua 0.82 0.51 0.77 0.20 5.0 Swift 0.85 0.52 0.72 0.23 6.5 Visual Basic 0.80 0.49 0.75 0.09 7.5 Cython 0.79 0.51 0.75 0.23 7.5 Julia 0.83 0.49 0.72 0.29 8.0 Moonscript 0.74 0.50 0.75 0.41 9.0 Sed 0.69 0.53 0.85 0.60 9.5 Scheme 0.78 0.49 0.71 0.14 11.5 Shell 0.74 0.49 0.74 0.39 12.0 Clojure 0.80 0.48 0.67 0.08 12.5 COBOL 0.73 0.49 0.74 0.03 12.5 F# 0.78 0.49 0.71 0.06 13.0 Octave 0.76 0.50 0.71 0.55 13.0 bc 0.73 0.50 0.67 0.55 16.5 Vim 0.69 0.49 0.67 0.48 16.5
Table 4: Score distribution for low resource target languages common for Clone Detection and Code Repair. Scores are the mean of the score of every source language for the given target language. Mean rank is used to rank the languages based on their ranking across each task. All languages benefit from transfer learning. Dart, TypeScript benefit the most and bc, Vim benefit the least.
@@
We do a similar analysis for low-resource target languages in clone detection and code repair, since these two tasks have a wider and more varied set of low-resource languages. Here, the most transferrable target languages from best to worst are Dart, TypeScript, Elixir, Lua, Swift, COBOL, F#, Octave, bc, and Vim. See Table 4 for details.
@@
For error detection, only a smaller set of languages can be considered low resource. Average scores vary less across different sources with highest being Rust with 0.66 and lowest being C# with 0.64. For solution domain classification there are no low-resource target languages.
While all target languages benefit from cross-lingual training, some benefit much more than others. Among low-resource target languages, Dart and TypeScript benefit the most.
Most Transferable Source Language.
Clone Detection Code Repair Solution Domain Error Detection Source score rank score rank score rank score rank Mean Rank Kotlin 0.8 3 0.78 3 0.5 1 0.67 1 2.0 JavaScript 0.82 1 0.81 1 0.47 4 0.62 6 3.0 Java 0.8 4 0.78 5 0.49 3 0.66 2 3.5 Python 0.73 6 0.79 2 0.49 2 0.66 5 3.75 C 0.8 2 0.78 4 0.45 5 0.66 4 3.75 C++ 0.79 5 0.78 6 0.44 6 0.66 3 5
Table 5: Source languages common for all tasks ranked by mean score across target languages and tasks. Score is the mean of the scores of all target languages for a given source language. We see that Kotlin is relatively the best source language and C++ the worst. This result is surprising since Kotlin is not a seen language during pre-training.
@@
The number of source languages vary for each of the four tasks. To understand the source language dependency and to identify the most transferrable source language, we start from languages all four tasks have in common (C, C++, Java, JavaScript, Kotlin, and Python). For each task, for each language, we calulate the mean cross-lingual score. Once we have the cross-lingual mean for each language, we rank the source languages based on their mean score for each task. Based on these rankings, we calculate the mean rank, for overall cross-lingual performance of source languages, across all target languages, across all tasks.
@@
The most transferrable source languages from best to worst are Kotlin, JavaScript, Java, Python, C, and C++. It may be surprising that Kotlin is the best-performing source language, since it is rarely discussed in the context of training LLMs and is not part of the pre-training corpus for CodeT5 (Wang et al., 2021). Prior work on transfer learning for natural languages (de Vries et al., 2022) yielded a similar surprise, with Romanian as the best-performing source language. A possible explanation could be that Romanian and Kotlin both have strong roots in dominant languages of the past (Latin and Java), while sitting at a cross-roads that helped them absorb many influences. Detailed scores and rankings are in Table 5.
Clone Detection Code Repair Error Detection
Source score rank score rank score rank Mean Rank
JavaScript 0.81 1 0.82 1 0.63 6 2.67
Java 0.79 2 0.75 5 0.66 1 2.67
Kotlin 0.78 3 0.77 3 0.66 4 3.33
C 0.77 4 0.76 4 0.66 2 3.33
Python 0.71 6 0.79 2 0.65 5 4.33
C++ 0.76 5 0.75 6 0.66 3 4.67
Table 6: Source languages common for all tasks ranked by mean score across low-resource target languages and three tasks. Solution Domain Classification is missing because all the target languages for this task are high-resource. Score is the mean of the scores of all target languages for a given source language. We see that JavaScript, Java are relatively the best source languages and C++ the worst.
@@
The above analysis considers all target languages, including high-resource target languages. In a more realistic setting, we can ignore high-resource target languages because the monolingual option is available for every high-resource language in every task. If we repeat the above analysis with only low-resource languages, we are left with a smaller set of target languages and only three tasks (all target languages for solution domain classification are high-resource). For the remaining three tasks and six common source languages, the most transferrable source languages from best to worst are JavaScript, Java, Kotlin, C, Python, and C++. JavaScript and Java contribute the most data to CodeT5 pretraining; Kotlin and C++ are absent (Wang et al., 2021). For details see Table 6.
Clone Detection Code Repair
Source score rank score rank Mean Rank
JavaScript 0.81 2 0.82 1 1.5
Scala 0.81 1 0.77 6 3.5
Haskell 0.81 4 0.78 3 3.5
Kotlin 0.78 11 0.77 5 8
D 0.81 3 0.71 15 9
Java 0.79 9 0.75 9 9
PHP 0.79 7 0.74 13 10
Nim 0.80 5 0.71 16 10.5
Python 0.71 19 0.79 2 10.5
Rust 0.79 8 0.72 14 11
Perl 0.79 6 0.68 17 11.5
C 0.76 15 0.76 8 11.5
Go 0.76 16 0.77 7 11.5
C# 0.77 12 0.74 12 12
bf 0.55 20 0.77 4 12
Fortran 0.78 10 0.67 18 14
Awk 0.76 17 0.75 11 14
C++ 0.76 18 0.75 10 14
Lisp 0.77 13 0.67 19 16
Pascal 0.77 14 0.66 20 17
Table 7: Source languages common for Clone Detection and Code Repair ranked by mean score across low-resource target languages. Score is the mean of the scores of all target languages for a given source language. We see that JavaScript, Scala, and Haskell are relatively the best source languages and C++, Lisp, and Pascal the worst.
@@
If we consider only clone detection and code repair, we get a much broader set of about 20 common source languages. Repeating the analysis for these two tasks for low-resource target languages only, we find once again that JavaScript is the most transferable and C++ is among the least. More details are in Table 7.
Kotlin is the best source language over all target languages and tasks. JavaScript is the best source language for low-resource target languages. C++ is a poor source language.
@@
Clone Detection Code Repair Error Detection
Target Source score Source score Source score
Crystal Ruby 0.93
Dart Scala 0.93 D 0.88
Swift JavaScript 0.91 Kotlin 0.79
bc Scala 0.91 JavaScript 0.87
TypeScript JavaScript 0.90 JavaScript 0.95
Elixer C 0.89 JavaScript 0.83
Clojure L 0.89 bf 0.77
Visual Basic Scala 0.89 JavaScript 0.81
Julia Ruby 0.89 JavaScript 0.81
Cython Python 0.88 Python 0.95
Lua Ruby 0.88 JavaScript 0.87
Moonscript Scala 0.87 JavaScript 0.87
Scheme Lisp 0.86 L 0.83
COBOL Haskell 0.85 Python 0.78
F# Haskell 0.84 Haskell 0.80
Sed perl 0.81 C 0.89
Octave ml 0.81 Python 0.78
Bash perl 0.80 bf 0.83
dc perl 0.80
Awk Vim 0.78
C# C++ 0.68
Rust Python 0.67
Ruby Kotlin, Java 0.67
PHP Kotlin, C++, C 0.67
Go Kotlin 0.67
Table 8: The best performing source language for each low resource target language, in each task. For high resource languages, the best performing source language is usually the same language. Ruby is high-resource for Clone Detection and not a target for Code Repair, hence its results are not show for these two tasks. Similarly C#, Rust, PHP, and Go are high-resource for Clone Detection and Code Repair.
@@
For clone detection in Figure 5(a), a model finetuned on Ruby does almost as well on Crystal, a language whose syntax is inspired by Ruby. Figure 5(a-b) also shows that surprisingly, Haskell does very well on COBOL, much better than FORTRAN does on COBOL. Lisp does very well on Scheme, which can be expected as Scheme is a dialect of Lisp. Similary Python does very well on Cython. Figure 5(a-d) also show that C-category languages (C, C++, and C#) do well on each other. Details are in Table 8.
Some language pairs show particularly good results. We need to analyze their features to understand why, motivating Section 4.3.
Dendrograms.
Dendrograms for each task cluster programming languages with similar transfer learning performances. In the code repair task, syntaxically and semantically related languages are next to each other. For example, C is paired with C++, JavaScript with TypeScript, Python with Cython, Lisp with Scheme. On the other hand, relatedness of languages plays a smaller role in the dendrograms for other tasks. This might be because while code repair requires deep language understanding, other tasks depend more upon other features. For instance, Ahmed & Devanbu (2022) show that some tasks depend heavily on identifiers.
Relatedness of languages plays a big role for code repair but less so for other tasks.
4.2 Performance Prediction
Given a task and target language, how should we pick the source language for best performance? Knowing this without having to train models on different source languages can save significant resources. It also provides a basis to study important features that characterize a successful transfer from a language to another. To predict the performance of the models on the different tasks for different source languages, we train a ranking model based on gradient-boosted decision trees (Ke et al., 2017).
We consider the features defined in Table 9, grouped into four categories. (1) Linguistic features: general properties characterizing a language pair like whether they support the same paradigms, the same style of memory management (garbage collector or not), or the same kind of type-system. (2) Syntactic features: properties of the syntax of a language pair measured by overlap in the counts of certain token types, as determined using the Pygments library on the corpus of code in the given languages. (3) Dataset-specific features: properties of the problems associated with code samples for the language pair in the online coding judge website from which the dataset came. (4) Model-specific features: whether the model saw the source, the target, or both languages during pretraining. The linguistic features of each language are given Table LABEL:tab:linguistic-features.